OpenStudy (joshiscool):

Whats the mean in this set of numbers {–11, 22, –2, –3}

OpenStudy (joshiscool):

gtg

TheSmartOne (thesmartone):

Add all the numbers and divide by how many numbers there are

OpenStudy (wolfgirl12):

the answer is 1.5 -11+22+(-3)+(-2)=6/4=1.5